# The Formality of the Goldman–Turaev Lie Bialgebra on a Closed Surface

### v1

@@ 4.3 Integration and Connections
Recall that the space of de Rham forms $\operatorname{DR}_{R}^{\bullet}\!B$ is the graded cyclic quotient of $\Omega^{\bullet}_{R}B$.
@@
Let $B$ and $M$ be equipped with a topology, $\mathfrak{d}$ a pro-unipotent Lie algebra and $(\mathfrak{d},\varphi,\rho)$ a continuous derivation action $M$. Let $\nabla$ be a trace-flat $R$-linear continuous connection on $M$, namely, $\operatorname{Tr}(\nabla^{2})=0$ in $\operatorname{DR}_{R}^{2}\!B$. We denote the integration of $\operatorname{Div}^{\nabla}\colon\mathfrak{d}\to|B|$ by $\mathsf{J}^{\nabla}\colon\operatorname{exp}(\mathfrak{d})\to|B|$. Then, for $G\in\operatorname{exp}(\mathfrak{d})$, we have $d(\mathsf{J}^{\nabla}(G))=\operatorname{Tr}((\operatorname{Ad}_{G}-\operatorname{id})\nabla)$ in $\operatorname{DR}_{R}^{1}\!B$.
@@ 5. Rewriting Rules
Let $T(H)_{\omega}=T(H)/\langle\omega\rangle$, which is the quotient of (non-completed) $T(H)$ by the two-sided ideal $\langle\omega\rangle$. The purpose of this section is to construct a basis of $|T(H)_{\omega}|$ for the later sections. See Theorem 5.22 for the final result.
@@ 6. The Closed Surface Case
In this section, we solve the formality problem for a closed surface $\Sigma=\Sigma_{g,0}$. The following is already known.
@@
Proof. This is a consequence of the formality on a surface with only one boundary component. Since a framing does not matter in this case, the set of formality isomorphisms is non-empty by Theorem 7.1 of [AKKN23] for any $g\geq 1$.∎
@@
We define connections on each piece:
@@
Proof. We have $F=\exp(f)$ for some $f\in\operatorname{Der}^{+}(\hat{L}(H)_{\omega})$. We show that any element $\operatorname{Der}^{+}(\hat{L}(H)_{\omega})$ can be lifted to $\operatorname{Der}_{\omega}^{+}(\hat{L}(H))=\{\tilde{f}\in\operatorname{Der}^{+}(\hat{L}(H)):\tilde{f}(\omega)=0\}$. Assume $f$ is homogeneous and take any lift $\tilde{f}$ of $f$ to $\operatorname{Der}^{+}(\hat{L}(H))$. We are done if $\tilde{f}(\omega)=0$, but it is not the case in general; instead, $\tilde{f}(\omega)$ is an element of the ideal $[H,[H,\cdots[H,\omega]\cdots]]$ generated by $\omega$. Since $\deg(f(\omega))\geq 1+\deg(\omega)=3$, we can write $\tilde{f}(\omega)=\sum_{i}[x_{i},a_{i}]+[y_{i},b_{i}]$ with $a_{i},b_{i}\in[H,[H,\cdots[H,\omega]\cdots]]$. Now define $f^{\prime}\in\operatorname{Der}^{+}(\hat{L}(H))$ by $f^{\prime}(x_{i})=\tilde{f}(x_{i})+b_{i}$ and $f^{\prime}(y_{i})=\tilde{f}(y_{i})-a_{i}$. Then, $f^{\prime}$ is also a lift of $f$, and
@@
Now we have $\operatorname{Div}^{\nabla_{\bullet,H}}(g)=|\tilde{\Delta}_{\omega}|(\operatorname{div}^{\nabla^{\prime}_{\bullet,H}}(g))$ by Section 6 of [Tan24b]. Since the antipode is bijective, it is equivalent to $|\Delta_{\omega}|(\operatorname{div}^{\nabla^{\prime}_{\bullet,H}}(g))=0\in\mathfrak{g}^{\otimes 2}$. In addition, $\operatorname{div}^{\nabla^{\prime}_{\bullet,H}}(g)$ has no constant term since the connection and hence the divergence both have degree 0 while $g$ has a positive degree. Therefore, using a canonical map $\mathfrak{g}\xrightarrow{\cong}|\hat{T}(H)_{\omega}|^{(\geq 1)}\subset|\hat{T}(H)_{\omega}|$, the equality is restated as

### v2

@@ 4.3 Integration and Connections
Recall that the space of de Rham forms $\operatorname{DR}_{R}^{\bullet}\!B$ is the graded trace space of $\Omega^{\bullet}_{R}B$.
@@
Let $B$ and $M$ be equipped with a topology, $\mathfrak{d}$ a pro-nilpotent Lie algebra and $(\mathfrak{d},\varphi,\rho)$ a continuous derivation action $M$. Let $\nabla$ be a trace-flat $R$-linear continuous connection on $M$, namely, $\operatorname{Tr}(\nabla^{2})=0$ in $\operatorname{DR}_{R}^{2}\!B$. We denote the integration of $\operatorname{Div}^{\nabla}\colon\mathfrak{d}\to|B|$ by $\mathsf{J}^{\nabla}\colon\operatorname{exp}(\mathfrak{d})\to|B|$. Then, for $G\in\operatorname{exp}(\mathfrak{d})$, we have $d(\mathsf{J}^{\nabla}(G))=\operatorname{Tr}((\operatorname{Ad}_{G}-\operatorname{id})\nabla)$ in $\operatorname{DR}_{R}^{1}\!B$.
@@ 5. Rewriting Rules
Consider the case $\Sigma=\Sigma_{g,0}$. The first homology $H=H_{1}(\Sigma;\mathbb{K})$ is a symplectic space and we have $\omega=\sum_{1\leq i\leq g}[x_{i},y_{i}]$. We put $T(H)_{\omega}=T(H)/\langle\omega\rangle$, which is the quotient of (non-completed) $T(H)$ by the two-sided ideal $\langle\omega\rangle$. The purpose of this section is to construct a basis of $|T(H)_{\omega}|$ for the later sections. See Theorem 5.22 for the final result.
@@ 6. The Closed Surface Case
In this section, we solve the formality problem for a closed surface $\Sigma=\Sigma_{g,0}$. The following is already known in [AKKN23].
@@
Proof. This is a consequence of the formality on a surface with only one boundary component. First of all, since a framing does not matter in this case, the set of formality isomorphisms is non-empty by Theorem 7.1 of [AKKN23] for any $g\geq 1$. Therefore, we can take a formality isomorphism $F\colon\widehat{\mathbb{K}\pi_{1}(\Sigma_{g,1})}\to\mathrm{gr}(\widehat{\mathbb{K}\pi_{1}(\Sigma_{g,1})})$. This descends to $F\colon\widehat{\mathbb{K}\pi_{1}(\Sigma_{g,0})}\to\mathrm{gr}(\widehat{\mathbb{K}\pi_{1}(\Sigma_{g,0})})$ as being the formality isomorphism forces $F$ to satisfy the condition $F(\xi)=\omega$. Since the Lie bialgebra structure of $|\widehat{\mathbb{K}\pi_{1}(\Sigma_{g,0})}/\mathbb{K}1|$ is the natural quotient of $|\widehat{\mathbb{K}\pi_{1}(\Sigma_{g,1})}|$, the claim follows.∎
@@
The proof of exactness can be obtained by taking the associated graded of the proof of Lemma 8.4 in [Tan24b]. Now we define connections on each piece:
@@
Proof. We have $F=\exp(f)$ for some $f\in\operatorname{Der}^{+}(\hat{L}(H)_{\omega})$. We will show that any element $\operatorname{Der}^{+}(\hat{L}(H)_{\omega})$ can be lifted to $\operatorname{Der}_{\omega}^{+}(\hat{L}(H))=\{\tilde{f}\in\operatorname{Der}^{+}(\hat{L}(H)):\tilde{f}(\omega)=0\}$. Assume $f$ is homogeneous and take any lift $\tilde{f}$ of $f$ to $\operatorname{Der}^{+}(\hat{L}(H))$. We are done if $\tilde{f}(\omega)=0$, but it is not the case in general; instead, $\tilde{f}(\omega)$ is an element of the ideal $[H,[H,\cdots[H,\omega]\cdots]]$ generated by $\omega$. Since $\deg(f(\omega))\geq 1+\deg(\omega)=3$, we can write $\tilde{f}(\omega)=\sum_{i}[x_{i},a_{i}]+[y_{i},b_{i}]$ with $a_{i},b_{i}\in[H,[H,\cdots[H,\omega]\cdots]]$. Now define $f^{\prime}\in\operatorname{Der}^{+}(\hat{L}(H))$ by $f^{\prime}(x_{i})=\tilde{f}(x_{i})+b_{i}$ and $f^{\prime}(y_{i})=\tilde{f}(y_{i})-a_{i}$. Then, $f^{\prime}$ is also a lift of $f$, and
@@
Now we have $\operatorname{Div}^{\nabla_{\bullet,H}}(g)=|\tilde{\Delta}_{\omega}|(\operatorname{div}^{\nabla^{\prime}_{\bullet,H}}(g))$ by Section 6 of [Tan24b]. Since the antipode is bijective, $\operatorname{Div}^{\nabla_{\bullet,H}}(g)=0$ is equivalent to $|\Delta_{\omega}|(\operatorname{div}^{\nabla^{\prime}_{\bullet,H}}(g))=0\in\mathfrak{g}^{\otimes 2}$. In addition, $\operatorname{div}^{\nabla^{\prime}_{\bullet,H}}(g)$ has no constant term since the connection and hence the divergence both have degree 0 while $g$ has a positive degree. Therefore, using a canonical map $\mathfrak{g}\xrightarrow{\cong}|\hat{T}(H)_{\omega}|^{(\geq 1)}\subset|\hat{T}(H)_{\omega}|$, the equality is restated as